OpenStudy (anonymous):

Someone please help :( The area of a rectangular piece of land is 260 square meters. If the length of the land was 5 meters less and the width was 2 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

really need you're help asap :( @Nnesha

OpenStudy (anonymous):

The rectangular area: xy = 260 The square area: (x + 2)(y - 5) = 260 Now, the thing about this is, if the area is square, then x = y, so (x + 2)(x - 5) = 260 x^2 - 3x - 10 = 260 x^2 - 3x - 270 = 0 (x - 15)(x + 18) = 0 x - 15 = 0 x + 18 = 0 Remember, only the positive result matters.

OpenStudy (anonymous):

What would be the answer to B tho...

OpenStudy (anonymous):

15

OpenStudy (anonymous):

and an equation I can use?

OpenStudy (anonymous):

(x + 2)(x - 5) = 260